# On the jump of the cover time in random geometric graphs

### v2

@@ 5.2. Proof of the upper bound on the cover time with fixed radius
Given any tessellation, let $C_{d}$ be the dimension-dependent constant so that every $d$-dimensional cube is adjacent (sharing any lower-dimensional face) to at most $C_{d}$ other cubes, and let $r_{0}=r_{0}(d)$ be the smallest radius $r$ so that $C_{d}e^{-V_{d}r^{d}}\leq 0.99$. Note that $r_{0}=\Theta(1)$. In the proof of Theorem 5.1 we showed that the effective resistance between any two vertices of the giant component of $G\in\mathcal{G}(n,r,d)$ is a.a.s. $O(\log n)$ for every $(1+\varepsilon)r_{g}\leq r\leq r_{0}$ (recall that the constant $\gamma$ in the inequality of the Theorem 5.1 is uniformly bounded from below for our range of $r$). Consider thus $r>r_{0}$. Observe that the graph $G\in\mathcal{G}(n,r_{0},d)$ can be naturally coupled with the graph $G^{\prime}\in\mathcal{G}(n,r,d)$ such that the giant of $G$ is a a subgraph of the giant of $G^{\prime}$. Hence, by Theorem 2.3, for vertices $x,y$ in both giants, the effective resistance between any two such vertices is also in $G^{\prime}$ a.a.s. bounded by $O(\log n)$.
By the triangle inequality for the effective resistance (Lemma 2.4), it thus remains to show that a.a.s. for any vertex $x$ that is in the giant of $G^{\prime}$ but not in the giant of $G$, there exists $x^{\prime}$ in the giant of $G$, such that $d_{G}(x,x^{\prime})=O(\log n)$.
To see this, note that in $G$, there is a.a.s. no diagonally connected set of cells of side length $r/\mu$ with integer left bottom corner inside $\Lambda_{n}$ of size $C\log n$ for $C$ large enough, with the property that all its cells are empty: indeed, letting $c:=C_{d}e^{-V_{d}r_{g}^{d}}\leq 0.99$, the expected number of such sets is at most $nc^{C\log n}$, since there are at most $n$ choices for the starting cell, at most $C_{d}$ choices for continuing at each step, and $e^{-V_{d}r_{g}^{d}}$ is the probability that the selected cell is indeed empty. Note that for $C=C(c)$ large enough, $nc^{C\log n}=o(1)$, and hence, by Markov’s inequality, a.a.s. such a set does not exist. Since every connected component different from the giant in $G$ has to have an empty diagonally connected outer boundary (if needed, we restrict to the outer boundary’s largest connected piece inside $\Lambda_{n}$, which is still of the same logarithmic order, only by a constant only depending on the dimension smaller than the total boundary), each such component thus has to have Euclidean boundary at most $O(\log n)$, and thus, for every $x^{\prime}$ in the giant of $G^{\prime}$, there must exist a vertex $x$ in the giant of $G$, such that $d_{E}(x,x^{\prime})=O(\log n)$. Then, by Corollary 2.12, a.a.s. also $d_{G}(x,x^{\prime})=O(\log n)$, and the theorem follows. ∎
Proof of Theorem 1.1. By combining Theorem 4.2 and Theorem 5.2, Theorem 1.1 follows immediately. ∎

### v3

@@ 6.1. Resistance bound in the backbone of dense cubes
First, make $C$ big enough such that $\beta r^{d}$ is at least 3 and that the dense cubes dominate a supercritical Bernoulli site percolation process. Divide $\Lambda_{n}$ into disjoint $\alpha$-logarithmic strips of cubes $\mathrm{Q_{i}}$. We use the structure introduced in Subsection 5.1, with the difference that we consider cubes $\mathrm{Q_{i}}$ instead of cubes $\mathrm{C_{i}}$, and dense cubes play the role of good cubes. Recall that we obtain the construction given there a.a.s.
In the following we assume vertices $x,y$ belonging to a fixed $2$-dimensional slice (we generalize this later). Let $\mathcal{Q}^{Y}_{X}$ (here $Y$ is a uniform discrete random variable independent from $X$ with values in $\{1,\dots,\lfloor c\log n\rfloor\}$, see the construction of Subsection 5.1 for the definition of $c$ and Theorem 5.1 for details of the construction of $\mathcal{Q}^{\ell}_{X}$ with a fixed value $\ell$ of $Y$) be the random path of dense cubes between two fixed dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, such that both belong to the giant component of dense cubes in $\mathcal{C}^{site}_{\mathrm{Q}}(n)$, and as in Subsection 5.1, we may assume that the two dense cubes belong to some parallelepipeds $\mathrm{Q_{i}}\in\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{Q_{j}}\in\mathrm{T}_{s_{2},r_{2}}$ such that $r_{1}=r_{2}$, $s_{2}>s_{1}$, $s_{2}-s_{1}$ is odd, and that both parallelepipeds do not belong to a strip that is at the boundary of the 2-dimensional slice. Pick vertices $x,y$ of $G\in\mathcal{G}(n,r,d)$ such that $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$. For each dense cube that is part of some realization of $\mathcal{Q}^{Y}_{X}$ select $\lfloor\beta r^{d}\rfloor$ vertices, except for $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$ where we may choose $\lfloor\beta r^{2}\rfloor-1$ vertices distinct from $x,y$ (or $\lfloor\beta r^{2}\rfloor-2$ in case $\mathrm{Q_{i}}=\mathrm{Q_{j}}$).
Conditional on $X$ and $Y$, we construct a random path of $G\in\mathcal{G}(n,r,d)$ that starts at $x$, finishes at $y$, and is entirely contained within the path of dense cubes $\mathcal{Q}_{X}^{Y}$ as follows. First, select uniformly at random and independently from $X,Y$ one of the $\lfloor\beta r^{d}\rfloor-1$ vertices that belong to the dense cube $\mathrm{Q_{i}}$ (one out of the $\lfloor\beta r^{d}\rfloor-1$ vertices previously selected), and connect it by an edge to the vertex $x$: this is the first edge of our path. For subsequent steps we do something similar: suppose that the last vertex that we chose belongs to the dense cube $\mathrm{Q}_{s}$. We then choose uniformly and independently at random from the $\lfloor\beta r^{d}\rfloor$ vertices that belong to the dense cube that follows in the path $\mathcal{Q}^{Y}_{X}$. We proceed until the last selected vertex belongs to the cube just before $\mathrm{Q_{j}}$ (the latter is thee dense cube that contains $y$). In this case we select uniformly at random between the $\lfloor\beta r^{d}\rfloor-1$ selected vertices in $\mathrm{Q_{j}}$ and finally connect all these last vertices by an edge to $y$ (if we had $\mathrm{Q_{i}}=\mathrm{Q_{j}}$, then we would pick $\lfloor\beta r^{d}\rfloor-2$ vertices inside the cube and form paths of length $2$ from $x$ to $y$ through these vertices). This process gives rise to $\lfloor\beta r^{d}\rfloor-2$ internally vertex-disjoint random paths of $G\in\mathcal{G}(n,r,d)$ with initial vertex $x$ and final vertex $y$. We denote this random path of $G\in\mathcal{G}(n,r,d)$ by $\mathcal{J}_{X,Z}^{Y}$ and denote its law by $\mathbb{P}_{X,Y,Z}$, where $Z$ denotes the randomness that comes from the path selection conditioned on $X,Y$. Observe that the number of edges that have some endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $O(r^{2d}\log n)$, since $\mathrm{V}^{\prime}(\mathrm{Q_{i}})$ (see also Subsection  5.1) is made of $\Theta(\log n)$ dense cubes, each of which has $O(r^{d})$ selected vertices, and any such vertex has $\Theta(r^{d})$ neighbors among selected vertices inside its own cube and inside a constant number of cubes around it. Since between any two consecutive dense cubes there are $\Theta(r^{2d})$ edges, and we choose uniformly at random one of them, we obtain in the case $u\neq x,u\neq y,v\neq x,v\neq y$,
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{Z}^{Y}|X,Y)=O\left(\frac{1}{r^{2d}}\right),$
whereas in the case |$|\{u,v\}\cap\{x,y\}|=1$, we have
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{X,Z}^{Y}|X,Y)=O\left(\frac{1}{r^{d}}\right).$
$+\sum_{t}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}p^{2}(u,v).$ (6.1)
To bound the first two sums, since $p(x,v)=O\left(\frac{1}{r^{d}}\right)$, and since $x$ and $y$ send flow to different $O(r^{d})$ vertices, we have $\sum_{v\in L_{1}}\big{(}p^{2}(y,v)+p^{2}(x,v)\big{)}=O\left(1/r^{d}\right)$. For the sum in the second row, since the number of edges with at least one endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $O(r^{2d}\log n)$, and since $p(u,v)=O(1/r^{2d})$,
$\sum_{\begin{subarray}{c}u\in\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\\ \end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\neq x,v\neq y\\ u\sim v\end{subarray}}p^{2}(u,v)=O\left(\frac{\log n}{r^{2d}}\right).$
To bound the term in the last row, recall that, as in the proof of Theorem 5.1 in Section 5, the probability that a dense cube inside of some parallelepiped $\mathrm{T}\in D_{t}$ belongs to $\mathcal{Q}_{X}^{Y}$ is $O\left(\frac{1}{t\log n}\right)$ for $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$ and $O\left(\frac{1}{\log n}\right)$ for $t=0$ (if the parallelepiped does not belong to $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$), since there are $\Omega(t)$ parallelepipeds at level $t\geq 1$, and we choose one of the $\Omega(\log n)$ crossings of dense cubes, uniformly at random. Now, observe that we have the additional randomness coming from the fact for any pair of consecutive dense cubes, we choose u.a.r. a vertex inside each cube, giving another factor $O(1/r^{2d})$, as mentioned above. All three choices are independent of each other, and thus, for edges $\{u,v\}$ appearing in the last term, we have
$p(u,v)=O\left(\frac{1}{tr^{2d}\log n}\right).$
The number of edges with positive probability $p$ and with at least one endpoint inside of a parallelepiped $\mathrm{T}$ is a.a.s. $O(r^{2d}\log^{2}n)$, as there are $O(\log^{2}n)$ dense cubes inside $\mathrm{T}$ (the volume of $\mathrm{T}$ is $\Theta(r^{d}\log^{2}n)$, and each cube has volume $\Theta(r^{d})$), each dense cube contains $O(r^{d})$ vertices, and the selected vertices are connected by an edge to another $O(r^{d})$ selected vertices by construction. Thus
$\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}$ $\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}p^{2}(u,v)$ (6.2)
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\\ v\neq x,v\neq y\end{subarray}}\frac{c^{\prime\prime\prime}}{t^{2}r^{4d}\log^{2}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}r^{2d}\log^{2}n\frac{c^{\prime\prime\prime\prime}}{t^{2}r^{4d}\log^{2}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}(t+1)\frac{c^{\prime\prime\prime\prime}}{t^{2}r^{2d}}$
$=O\left(\frac{1}{r^{2d}}\right)\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}O\left(\frac{1}{t}\right)$
$=O\left(\frac{\log n}{r^{2d}}\right).$
It is clear that the contribution of $D_{0}$ to the formula is $O\left(\frac{1}{r^{2d}}\right)$. By symmetry, the contribution of $D_{t}$ with $\frac{s_{2}-s_{1}}{2}<t\leq s_{2}-s_{1}+1$ is also $O\left(\frac{1}{tr^{2d}}\right)$, and thus
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right).$
Recall that we assumed that $x,y$ belong to some dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, that at the same time are inside of some parallelepipeds in the same horizontal strip and separated by an even number of parallelepipeds. Before we lift these conditions, we first provide a useful construction: we claim that if $x^{\prime},y^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{Q_{i}}^{\prime}$ and $y^{\prime}\in\mathrm{Q_{j}}^{\prime}$ with both cubes being dense, and such that $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$ are connected by a path $\mathcal{L}$ of dense cubes of length $O(\log n)$, then $\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=\left(\frac{\log n}{r^{2d}}\right)$.
To see this, let $x^{\prime}$ send an amount of $\frac{1}{\lfloor\beta r^{d}\rfloor-1}$ flow to the $\lfloor\beta r^{d}\rfloor-1$ previously selected vertices of $\mathrm{Q_{i}}^{\prime}$. From there on, let the previously selected vertices to send a uniformly distributed flow to the selected $\lfloor\beta r^{d}\rfloor$ vertices that belong to the dense cube that follows in the path that connects $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$. We continue distributing in a uniform way the flow until we arrive until $\mathrm{Q_{j}}^{\prime}$. Once the flow reached the $\lfloor\beta r^{d}\rfloor-1$ selected vertices of $\mathrm{Q_{j}}^{\prime}$, send all the flow back to $y^{\prime}$. Denote this flow by $\theta$. We then have
As before, the first two sums contribute $O(r^{d})O(1/r^{2d})=O(1/r^{d})$. For the third sum, note that there are $O(\log n)$ dense cubes in $\mathcal{L}$, and between two consecutive dense cubes there are $O(r^{2d})$ edges with flow $O\left(\frac{1}{r^{2d}}\right)$ on each of them. Thus
$\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=O\left(\frac{1}{r^{d}}\right)+\sum_{|\mathcal{L}|}O\left(r^{2d}\right)O\left(\frac{1}{r^{4d}}\right)=O\left(\frac{\log n}{r^{2d}}\right),$
and the claim is proved.
We now show how this claim allows us to prove the bound on the effective resistance for any $x\in\mathrm{C_{i}}^{\prime}\in\mathrm{T}_{s_{1},r_{1}}$ and $y\in\mathrm{C_{j}}^{\prime}\in\mathrm{T}_{s_{2},r_{2}}$ in the same strip, first assuming that both parallelepipeds do not belong to strips on the boundary. First, if $x$ and $y$ are separated by an odd number of parallelepipeds (that is, $s_{2}-s_{1}$ even), we find $x^{\prime}$ in a parallelepiped adjacent to $x$, at Euclidean distance (in the renormalized lattice of cubes) $O(\log n)$ from $x$. Thus, by the claim, $\mathcal{R}(x\longleftrightarrow x^{\prime})=O\left(\frac{\log n}{r^{2d}}\right)$, and since $x^{\prime}$ and $y$ are separated by an even number of parallelepipeds, by the above, $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right)$, and finally by Lemma 2.4, also $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right)$. As before, by the same idea, we may extend our bounds also to arbitrary vertices inside the same parallelepiped (in cubes not satisfying all restrictions) and to vertices in parallelepipeds of a strip on the boundary.
To generalize the inequality between any pair of strips we use the same idea as in Section 5 from Figure 12: for any pair of strips, we find another vertex $z$ in the intersection of the strips that contain $x$ and $y$ and that intersect at some parallelepiped $\mathrm{T}$ that contains the vertex $z$ such that $\mathcal{R}(x\longleftrightarrow z)$ and $\mathcal{R}(y\longleftrightarrow z)$ are $O\left(\frac{\log n}{r^{2d}}\right)$ (such $z$ must exist, as there is a logarithmic amount of crossings of dense cubes in each strip), we apply again the triangle inequality and we are done.
To generalize to different slices, we proceed as in the last step of the proof of Theorem 5.1, were we saw that we may go from one slice to another in a (fixed) maximum of steps. Subsequent applications of the triangle inequality are enough to obtain the desired result.
We thus finished proving that a.a.s. the effective resistance between any two vertices that belong to dense cubes that belong to the giant component of dense cubes is $O\left(\frac{\log n}{r^{2d}}\right)$. ∎
Lemma 6.2.
Let $r_{0}$ be a large enough constant, and let $r_{0}\leq r\leq(1-\varepsilon)r_{c}$. Then
$\tau_{cov}(L_{1}(G))=O(n\log^{2}n)$ (6.4)
Proof.
It is a standard topological fact that every (maximal) connected component of cubes, which might contain dense and non-dense cubes and that has empty intersection with the giant component of dense cubes, has a diagonally connected outer boundary of non-dense cubes. As there are at most $(3^{d}-1)^{L}$ diagonally connected paths of length $L$ when the starting vertex is fixed, and since the probability that a cube is non-dense is $e^{-\kappa r^{d}}$ for some $\kappa=\kappa(\beta)>0$, the probability that there exists a diagonally connected path of non-dense cubes of length greater or equal than $L$ is, by a union bound over all $(1+o(1))n$ starting vertices, at most
$(1+o(1))ne^{-L\kappa r^{d}}(3^{d}-1)^{L},$
which tends to 0 for $L:=c\frac{\log n}{r^{d}}$ with $c>0$ large enough. Therefore, since for any $M>0$, every connected component (in the renormalized lattice of cubes $\mathrm{Q}$) of volume $M^{d}$, by a standard isoperimetric inequality has to have an outer boundary of length at least changed this, it was only for 2 dimension $CM^{d-1}$ for some $C>0$, a.a.s. every component different from the giant of the associated Bernoulli site percolation has volume at most $O(L^{d/(d-1)})$. In other words, for every site, at distance $L$ one finds a site belonging to the giant component (of the renormalized lattice of dense cubes), and for any vertex $x$ of the underlying random geometric graph, at Euclidean distance at most $O(Lr/\mu)=O(\log n/r^{d-1})$ we find a vertex $y$ belonging to a cube of the giant component of the Bernoulli site percolation.
Then, by Corollary 2.12, a.a.s. we have $d_{G}(x,y)=O\left(\log n/r^{d}\right)$, and this implies that $\mathcal{R}(x\longleftrightarrow y)=O(\log n/r^{d})$. If $z\in G$ is another vertex that belongs to some dense cube in the backbone $\mathcal{C}^{site}_{\mathrm{Q}}(n)$ then by  Theorem 6.1 we have $\mathcal{R}(y\longleftrightarrow z)=O(\log n/r^{2d})$, and hence by the triangle inequality, also
$\mathcal{R}(x\longleftrightarrow z)=O(\log n/r^{d})+O(\log n/r^{2d})=O(\log n/r^{d}).$
Since by Lemma 2.10, the number of edges of the giant component satisfies $|E|=\Theta(nr^{d})$ a.a.s., by Theorem 2.6, we have that
$\tau_{cov}(L_{1}(G))=O(n\log^{2}n),$
and the lemma follows. ∎
Proof of Theorem 1.1. By combining Theorem 4.2 and Theorem 5.1 with Lemma 6.2, Theorem 1.1 follows immediately. ∎